a plane leaves an airport traveling at an average speed of 240 kilometres per hour. how long will it take a second plane traveling the same route at an average speed of 600 kilometers per hour to catch up with the first plane if it leaves 3 hours later?
Answer by checkley71(8403)

240X=600(X-3)
240X=600X-1800
240X-600X=-1800
-360X=-1800
X=-1800/-360
X=5 HOURS IS THE TIME FOR THE FASTER PLANE TO CATCH UP WITH THE SLOWER PLANE.
PROOF
240*5=600*2
1200=1200 KM TRAVELED.